# Low-degree mod 2 cohomology of classifying spaces of $G_{2}$-gauge groups

### v2

@@ 4.1. The $E_{2}$–page in low degrees
By Proposition 3.2, the fibre cohomology $M^{t}$ vanishes for $1\leq t\leq 4$. Consequently, in total degree $j=s+t\leq 10$ the only nonzero groups $E_{2}^{s,t}$ with $t>0$ occur when
$t\geq 5.$
In particular, the class $u_{5}\in M^{5}$ gives rise to a class
@@
and this is the first positive-degree generator of $M$.
@@ 4.2. Bidegree constraints on differentials
Let $u_{5}\in E_{2}^{0,5}(k)\cong M^{5}$ be a nonzero class representing the first positive-degree fibre cohomology. In total degree $j\leq 10$ all differentials $d_{r}(u_{5})$ vanish for $2\leq r\leq 5$ and for $r\geq 7$, and the only page on which $u_{5}$ can support a nontrivial differential is $r=6$. More precisely,
@@
In total degree $s+t\leq 10$ the following hold.
@@
    For any $r\geq 7$, every differential
    $d_{r}\colon E_{r}^{s,t}(k)\longrightarrow E_{r}^{s+r,t-r+1}(k)$
    has either source or target lying in total degree >10>10. In particular, no $d_{r}$ with $r\geq 7$ affects the groups $E_{r}^{s,t}(k)$ with $s+t\leq 10$.
@@
    If we restrict to the $H^{*}(BG_{2})$–submodule generated by $u_{5}\in E_{2}^{0,5}(k)$, then all differentials $d_{r}$ with $2\leq r\leq 5$ vanish in total degree $s+t\leq 10$, and every differential
    $d_{6}\colon E_{6}^{s,t}(k)\longrightarrow E_{6}^{s+6,t-5}(k)$
    in this range is determined multiplicatively by the value of $d_{6}$ on $u_{5}$. Concretely, for $s+t\leq 10$ and any $x\in H^{s}(BG_{2})$ we have
    $d_{6}(x\cdot u_{5})=x\cdot d_{6}(u_{5})\in E_{6}^{s+6,0}(k),$
    and there are no other possible nonzero differentials on this submodule. In particular, up to multiplication by elements of $H^{*}(BG_{2})$, the only potentially nonzero differential in the range $s+t\leq 10$ with source in this submodule is
    $d_{6}\colon E_{6}^{0,5}(k)\longrightarrow E_{6}^{6,0}(k).$
@@
We examine all possible differentials
$d_{r}\colon E_{r}^{s,t}(k)\longrightarrow E_{r}^{s+r,t-r+1}(k)$
with $s+t\leq 10$.
For part (a), suppose $r\geq 7$ and $E_{r}^{s,t}\neq 0$ with $s+t\leq 10$. Then the target bidegree (s+r,tr+1)(s+r,t-r+1) has total degree
$(s+r)+(t-r+1)=s+t+1\leq 11.$
If $s+t\leq 10$ and $r\geq 7$, a short check shows that as soon as $s+r\geq 11$ or tr+1<0t-r+1<0 the target lies either in total degree >10>10 or in negative fibre degree. In each case the target group is zero in our range of interest, so no such $d_{r}$ can hit or originate from classes with $s+t\leq 10$. This proves (a).
For part (b) we restrict to the $H^{*}(BG_{2})$–submodule generated by $u_{5}$, i.e. to classes of the form
@@
with $s+5\leq 10$. Since $H^{s}(BG_{2})$ is nonzero only for $s\in\{0,4,6,7,8,10,\dots\}$, the only possibilities with $s+5\leq 10$ are $s=0$ and $s=4$, giving bidegrees $(0,5)$ and $(4,5)$.
We first analyse the differentials on $u_{5}$ itself. By Lemma 4.1 we know that $d_{r}(u_{5})=0$ for $2\leq r\leq 5$ and for $r\geq 7$, and that the only page on which a nontrivial differential can occur is $r=6$, with
$d_{6}(u_{5})\in E_{6}^{6,0}(k)\cong H^{6}(BG_{2}).$
Now let $x\in H^{s}(BG_{2})$ with $s+5\leq 10$ and consider the class $x\cdot u_{5}\in E_{2}^{s,5}(k)$.
(i) The case $2\leq r\leq 5$. Any differential with source $x\cdot u_{5}$ has target in bidegree (s+r,5r+1)(s+r,5-r+1). For $2\leq r\leq 5$ we have $5-r+1\in\{4,3,2,1\}$, so by Proposition 3.2 the fibre term $M^{5-r+1}$ vanishes and hence $E_{2}^{s+r,5-r+1}(k)=0$. Thus $d_{r}(x\cdot u_{5})=0$ for all $2\leq r\leq 5$.
(ii) The case $r=6$. Here the target of $d_{6}$ is
@@
The Serre spectral sequence is multiplicative, so $d_{6}$ satisfies the Leibniz rule
@@
But $x$ comes from the base and lies in filtration degree $(s,0)$, so $d_{6}(x)=0$ for degree reasons (its target would have negative fibre degree). Hence
@@
as claimed.
In the present range $s+5\leq 10$ the only nontrivial possibility is $s=4$, yielding a class in bidegree $(4,5)$ and a target in bidegree $(10,0)$. This shows that, within the submodule generated by $u_{5}$, all $d_{r}$ with $2\leq r\leq 5$ vanish and every $d_{6}$ is obtained from $d_{6}(u_{5})$ by multiplication with elements of $H^{*}(BG_{2})$.
Combining this with part (a) completes the proof of the lemma. ∎
Remark 4.3.
  • (i)
    By Lemma 4.1, the class $u=1\otimes u_{5}\in E_{2}^{0,5}(k)$ survives at least to $E_{6}^{0,5}(k)$, and the only potentially nonzero differential with source $u$ in total degree $j\leq 10$ is
    $d_{6}(u)\in E_{6}^{6,0}(k)\cong H^{6}(BG_{2})\otimes M^{0}\cong H^{6}(BG_{2}).$
    Since $H^{6}(BG_{2})\cong\mathbb{F}_{2}\{x_{6}\}$, there exists a uniquely determined scalar $\varepsilon(k)\in\mathbb{F}_{2}$ such that
    $d_{6}(u)=\varepsilon(k)\,x_{6}.$
    This is the scalar appearing in Theorem 1.1(iii).
  • (ii)
    Lemma 4.2 only analyzes the differentials with source in the $H^{*}(BG_{2})$–submodule
    $U\;=\;H^{*}(BG_{2})\cdot u_{5}\;\subset\;E_{2}^{*,*}(k).$
    We do not claim that $M^{j}=H^{j}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})$ vanishes for $j\geq 6$, and additional fibre generators in these degrees could in principle support nontrivial differentials in the Serre spectral sequence. Such differentials, however, do not change the description of the possible $d_{6}$–differentials on $U$ in total degree $\leq 10$, which is all that is needed for Theorem 1.1. A determination of $M$ in higher degrees would require an explicit computation via Eilenberg–Moore spectral sequences, and lies beyond the scope of the present paper.
Note 4.4.
@@
where $\mathrm{Map}_{0}(S^{4},BG_{2})$ denotes the component containing the constant map. The evaluation map $\mathrm{ev}\colon\mathrm{Map}_{0}(S^{4},BG_{2})\to BG_{2}$ admits a section $s\colon BG_{2}\to\mathrm{Map}_{0}(S^{4},BG_{2})$ defined by sending $y\in BG_{2}$ to the constant map at $y$. The existence of this section implies that the homomorphism
@@
is injective. In the Serre spectral sequence for the evaluation fibration with $k=0$, the class $x_{6}\in H^{6}(BG_{2})\cong E_{2}^{6,0}$ must therefore survive to $E_{\infty}$ and cannot lie in the image of any differential. In particular, $x_{6}$ cannot be hit by $d_{6}(u_{5})$, forcing $d_{6}(u_{5})=0$. Thus, in our notation, $\varepsilon(0)=0$.
Lemma 4.5 (A 22–local periodicity of $\varepsilon(k)$).
Work 22–locally. If $k\equiv k^{\prime}\pmod{4}$, then
@@
In particular, $\varepsilon(k)$ depends only on the residue class of $k$ modulo $4$.
@@
By [8, Lemma 2.1], the connecting map $\partial_{k}\colon G_{2}\to\Omega^{3}_{0}G_{2}$ is the triple adjoint of the Samelson product $\langle k\cdot i_{3},1\rangle$, and in particular $\partial_{k}\simeq k\cdot\partial_{1}$. Kameko proved that the 22–primary component of $\langle i_{3},1\rangle$ has order $4$ [7, Theorem 1.1]. Equivalently, after localization at $2$ the map $\partial_{1}$ has order $4$, so $\partial_{k}\simeq\partial_{k^{\prime}}$ 22–locally whenever $k\equiv k^{\prime}\pmod{4}$.
The Serre spectral sequence (1.1) is functorial with respect to maps of fibrations. Since the kk–dependence of the evaluation fibration is controlled by $\partial_{k}$ (see Section 3), the homotopy $\partial_{k}\simeq\partial_{k^{\prime}}$ induces an identification of the Serre spectral sequences for $k$ and $k^{\prime}$ which is compatible with the standard identifications $E_{2}^{s,t}(k)\cong H^{s}(BG_{2})\otimes H^{t}(\Omega^{3}_{0}G_{2})$. Under this identification, the classes $u_{5}\in E_{2}^{0,5}$ and $x_{6}\in E_{2}^{6,0}\cong H^{6}(BG_{2})$ correspond, and hence the $d_{6}$–values correspond. Since $H^{6}(BG_{2})\cong\mathbb{F}_{2}\{x_{6}\}$ is one-dimensional, the scalar in
@@
must coincide for $k$ and $k^{\prime}$, proving $\varepsilon(k)=\varepsilon(k^{\prime})$. ∎
@@
Work 22–locally. If $4\mid k$, then $\varepsilon(k)=0$.
@@
If $4\mid k$ then $k\equiv 0\pmod{4}$, so Lemma 4.5 gives $\varepsilon(k)=\varepsilon(0)$. By Note 4.4 we have $\varepsilon(0)=0$, hence $\varepsilon(k)=0$. ∎
@@
Lemma 4.5 and Corollary 4.6 show that, in the low-degree range considered here, the potential kk–dependence through the first fibre class $u_{5}$ reduces to at most three cases:
$k\equiv 0,\ 2,\ 1\text{ or }3\pmod{4}$.
In particular, every class $k\equiv 0\pmod{4}$ falls into the same low-degree behaviour as the trivial bundle, namely $\varepsilon(k)=0$. Thus, to detect genuinely new behaviour via $d_{6}(u_{5})$, it suffices to determine $\varepsilon(1)$ and $\varepsilon(2)$. This reduction is consistent with the 22–local classification of $G_{2}$–gauge groups governed by the 22–primary order of $\langle i_{3},1\rangle$ [8, 7].
4.3. The $E_{\infty}$–page and completion of the proof
We now describe the $E_{\infty}$–page in total degree $\leq 10$. By Lemma 4.2, the only potentially nonzero differential with source or target in this range on the submodule generated by $u_{5}$ is $d_{6}$ on $E_{6}^{0,5}(k)$. All other classes in total degree $\leq 10$ are permanent cycles.
Let us first consider the case $\varepsilon(k)=0$. Then $d_{6}(u)=0$, and by multiplicativity $d_{6}=0$ on all of $E_{6}^{*,*}(k)$ in total degree $\leq 10$. It follows that $E_{2}^{*,*}(k)\cong E_{\infty}^{*,*}(k)$ in this range, and the Serre spectral sequence collapses at $E_{2}$ in total degree $\leq 10$.
If $\varepsilon(k)=1$, then $d_{6}(u)=x_{6}$, and the class $x_{6}$ does not survive to $E_{\infty}^{6,0}(k)$. More generally, the image of $d_{6}$ is the $H^{*}(BG_{2})$–submodule generated by $x_{6}$, and $E_{\infty}^{*,*}(k)^{\leq 10}$ is obtained from $H^{*}(BG_{2})\otimes M^{\leq 10}$ by imposing the single relation $x_{6}=d_{6}(u_{5})$ in total degree $\leq 10$.
In either case, for each total degree $j\leq 10$ there is a finite filtration of $H^{j}(B\mathcal{G}_{k})$ whose associated graded pieces are the groups $E_{\infty}^{s,t}(k)$ with $s+t=j$. This gives the short exact sequence of Theorem 1.1(v), with $Q_{j}(k)$ recording possible extension phenomena between different bidegrees. Since the only dependence on $k$ in the spectral sequence occurs via the scalar $\varepsilon(k)$, it follows that all potential kk–dependence in $H^{j}(B\mathcal{G}_{k})$ for $j\leq 10$ is encoded in this single differential.
Proof of Theorem 1.1.
Part (i) is the identification of the $E_{2}$–term with trivial local coefficients. Part (ii) is Proposition 3.2, and parts (iii) and (iv) follow from Lemmas 4.1 and 4.2, together with the discussion above of the differential $d_{6}(u)=\varepsilon(k)x_{6}$. Part (v) is the standard description of $H^{*}(B\mathcal{G}_{k})$ in terms of the $E_{\infty}$–term and the resulting finite filtration. This completes the proof. ∎
Remark 4.8.
The restriction to total degree $j\leq 10$ in Theorem 1.1 is motivated by two factors. First, this range is sufficient to capture the first nontrivial cohomological dependence on the bundle class $k$, which manifests via the differential $d_{6}(u_{5})$. Second, extending the analysis beyond degree $10$ requires a substantially more detailed understanding of the fibre cohomology $M=H^{*}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})$ as an unstable $\mathcal{A}$–module, which involves Eilenberg–Moore spectral sequence computations beyond the scope of this paper. Once $M$ is explicitly determined in a larger range, the computation of $H^{*}(B\mathcal{G}_{k})$ reduces to a finite problem in the “hit” theory over the polynomial algebra $H^{*}(BG_{2})$, well suited to computer implementation in systems such as SageMath. We hope to return to these global computations in future work.

### v4

@@ 4.1. The $E_{2}$–page in low degrees
By Proposition 3.2, the fibre cohomology $M^{t}$ vanishes for $1\leq t\leq 4$. Consequently, in total degree $j=s+t\leq 10$ the only nonzero groups $E_{2}^{s,t}$ with $t>0$ occur when $t\geq 5.$ In particular, the class $u_{5}\in M^{5}$ gives rise to a class
@@
and this is a generator of least positive degree in $M$.
@@ 4.2. Bidegree constraints on differentials
Let $u_{5}\in E_{2}^{0,5}(k)\cong M^{5}$ be a nonzero class representing the fibre cohomology class of least positive degree. In total degree $j\leq 10$ all differentials $d_{r}(u_{5})$ vanish for $2\leq r\leq 5$ and for $r\geq 7$, and the only page on which $u_{5}$ can support a nontrivial differential is $r=6$. More precisely,
@@
Let $U\subset E_{2}^{*,*}(k)$ denote the $H^{*}(BG_{2})$–submodule generated by $u_{5}\in E_{2}^{0,5}(k)$. For classes in $U$ of total degree at most $10$ (equivalently, for classes $x\cdot u_{5}\in E_{2}^{s,5}(k)$ with $s+5\leq 10$), the following hold.
@@
    All differentials $d_{r}$ with $2\leq r\leq 5$ vanish on $U$.
@@
    Every differential
    $d_{6}\colon E_{6}^{s,5}(k)\longrightarrow E_{6}^{s+6,0}(k)$
    with $s+5\leq 10$ satisfies
    $d_{6}(x\cdot u_{5})=x\cdot d_{6}(u_{5})\qquad\text{for }x\in H^{s}(BG_{2}).$
  3. (c)
    All differentials $d_{r}$ with $r\geq 7$ vanish on $U$.
@@
By definition,
$U=H^{*}(BG_{2})\cdot u_{5}\subseteq E_{2}^{*,*}(k),$
so every class in $U$ of total degree $\leq 10$ has the form
@@
with $s+5\leq 10$.
Since
$H^{*}(BG_{2})\cong\mathbb{F}_{2}[x_{4},x_{6},x_{7}],$
the only values of $s$ for which $H^{s}(BG_{2})\neq 0$ and $s+5\leq 10$ are $s=0$ and $s=4$. Thus, in the range under consideration, the only possible bidegrees in $U$ are $(0,5)$ and $(4,5)$.
Proof of (a). Let $x\cdot u_{5}\in U$ with $s+5\leq 10$, and let $2\leq r\leq 5$. Then
$d_{r}(x\cdot u_{5})\in E_{r}^{s+r,\,5-r+1}(k).$
For $2\leq r\leq 5$ we have
$5-r+1\in\{4,3,2,1\}.$
By Proposition 3.2,
$M^{j}=H^{j}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})=0\qquad\text{for }1\leq j\leq 4.$
Hence the fibre term in bidegree (s+r,5r+1)(s+r,5-r+1) vanishes, so
$E_{r}^{s+r,\,5-r+1}(k)=0.$
Therefore
$d_{r}(x\cdot u_{5})=0\qquad\text{for all }2\leq r\leq 5.$
This proves (a).
Proof of (b). Now let $r=6$. Then
@@
Since the Serre spectral sequence is multiplicative, the differential satisfies the Leibniz rule:
@@
But $x$ lies in bidegree $(s,0)$, so any differential on $x$ would land in bidegree (s+6,5)(s+6,-5), which is zero for degree reasons. Thus
$d_{6}(x)=0,$
and consequently
@@
as claimed. This proves (b).
Proof of (c). Let $r\geq 7$. For any class $x\cdot u_{5}\in U$ with $s+5\leq 10$, one has
$d_{r}(x\cdot u_{5})\in E_{r}^{s+r,\,5-r+1}(k).$
Since $r\geq 7$,
$5-r+1=6-r\leq-1.$
Thus the target has negative fibre degree, so
$E_{r}^{s+r,\,5-r+1}(k)=0.$
Hence
$d_{r}(x\cdot u_{5})=0\qquad\text{for all }r\geq 7.$
This proves (c). ∎
Note 4.3.
@@
where $\mathrm{Map}_{0}(S^{4},BG_{2})$ denotes the path component containing the constant map. The evaluation map $\mathrm{ev}\colon\mathrm{Map}_{0}(S^{4},BG_{2})\to BG_{2}$ admits a section $s\colon BG_{2}\to\mathrm{Map}_{0}(S^{4},BG_{2})$ defined by sending $y\in BG_{2}$ to the constant map at $y$. The existence of this section implies that the homomorphism
@@
is injective. Indeed, one has
$s^{*}\circ\mathrm{ev}^{*}=\mathrm{id}_{H^{*}(BG_{2})},$
so $\mathrm{ev}^{*}$ is split injective.
In the Serre spectral sequence for the evaluation fibration with $k=0$, the class $x_{6}\in H^{6}(BG_{2})\cong E_{2}^{6,0}$ must therefore survive to $E_{\infty}$ and cannot lie in the image of any differential. In particular, $x_{6}$ cannot be hit by $d_{6}(u_{5})$, forcing $d_{6}(u_{5})=0$. Thus, in our notation, $\varepsilon(0)=0$.
Lemma 4.4 (An 88–periodicity statement for $\varepsilon(k)$).
Work 22–locally. If $k\equiv k^{\prime}\pmod{8}$, then
@@
In particular, $\varepsilon(k)$ depends only on the residue class of $k$ modulo $8$.
@@
By [11, Lemma 2.1], the connecting map
$\partial_{k}\colon G_{2}\to\Omega^{3}_{0}G_{2}$
is the triple adjoint of the Samelson product $\langle k\cdot i_{3},1\rangle$, and in particular
$\partial_{k}\simeq k\cdot\partial_{1}.$
Kameko proved that the 22–primary component of $\langle i_{3},1\rangle$ has order $4$ [10, Proposition 1.3]. Equivalently, after localization at $2$ the map $\partial_{1}$ has order $8$, so
$\partial_{k}\simeq\partial_{k^{\prime}}\qquad\text{$2$--locally whenever }k\equiv k^{\prime}\pmod{8}$.
Now the evaluation fibration
is a delooping of the homotopy fibration
$\mathcal{G}_{k}\longrightarrow G_{2}\xrightarrow{\;\partial_{k}\;}\Omega^{3}_{0}G_{2}.$
Hence, if $\partial_{k}\simeq\partial_{k^{\prime}}$ 22–locally, then the two homotopy fibrations over $G_{2}$ have equivalent homotopy fibres, and after delooping we obtain a fibre homotopy equivalence between the corresponding evaluation fibrations over $BG_{2}$. In particular, there is an isomorphism between the associated mod $2$ Serre spectral sequences which is compatible with the standard identifications
$E_{2}^{s,t}(k)\cong H^{s}(BG_{2})\otimes H^{t}(\Omega^{3}_{0}G_{2}),\qquad E_{2}^{s,t}(k^{\prime})\cong H^{s}(BG_{2})\otimes H^{t}(\Omega^{3}_{0}G_{2}).$
Under this identification, the class
$u_{5}\in E_{2}^{0,5}\cong H^{5}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})$
corresponds to the same generator of the fibre cohomology, and
$x_{6}\in E_{2}^{6,0}\cong H^{6}(BG_{2};\mathbb{F}_{2})$
corresponds to the same degree-$6$ generator on the base. Therefore the $d_{6}$–differentials agree. Since $H^{6}(BG_{2};\mathbb{F}_{2})\cong\mathbb{F}_{2}\{x_{6}\}$ is one-dimensional, the scalar in
@@
must be the same for $k$ and $k^{\prime}$. Thus
$\varepsilon(k)=\varepsilon(k^{\prime}),$
as required. ∎
@@
Work 22–locally. If $8\mid k$, then $\varepsilon(k)=0$.
@@
If $8\mid k$ then $k\equiv 0\pmod{8}$, so Lemma 4.4 gives $\varepsilon(k)=\varepsilon(0)$. By Note 4.3 we have $\varepsilon(0)=0$, hence $\varepsilon(k)=0$. ∎
4.3. The $E_{\infty}$–page and completion of the proof
For each total degree $j\leq 10$, the Serre spectral sequence yields a finite filtration
$0=F^{j+1}H^{j}(B\mathcal{G}_{k})\subseteq F^{j}H^{j}(B\mathcal{G}_{k})\subseteq\cdots\subseteq F^{0}H^{j}(B\mathcal{G}_{k})=H^{j}(B\mathcal{G}_{k})$
whose associated graded object is
$\mathrm{gr}\,H^{j}(B\mathcal{G}_{k})\cong\bigoplus_{s+t=j}E_{\infty}^{s,t}(k).$
By Lemmas 4.1 and 4.2, the contribution of the $H^{*}(BG_{2})$–submodule
$U=H^{*}(BG_{2})\cdot u_{5}\subseteq E_{2}^{*,*}(k)$
to $E_{\infty}^{*,*}(k)$ in total degree $j\leq 10$ is completely determined by the single differential
$d_{6}(u_{5})=\varepsilon(k)x_{6}.$
More precisely, within $U$ all differentials $d_{r}$ for $2\leq r\leq 5$ vanish, every $d_{6}$ is obtained from $d_{6}(u_{5})$ by multiplication with classes from $H^{*}(BG_{2})$, and no $d_{r}$ with $r\geq 7$ contributes in total degree $\leq 10$.
We do not claim here that these are the only possible differentials in total degree $\leq 10$ on the whole spectral sequence, since additional fibre classes in degrees $t\geq 6$ may in principle support further differentials. Thus Theorem 1.1 should be interpreted as a precise determination of the first kk–dependent contribution coming from the fibre class $u_{5}$.
Proof of Theorem 1.1.
Part (i) is the identification of the $E_{2}$–term with trivial local coefficients. Part (ii) is Proposition 3.2. Parts (iii) and (iv) follow from Lemmas 4.1 and 4.2. Part (v) is the standard filtration statement for a convergent cohomology Serre spectral sequence, together with the preceding description of the contribution of the submodule $U=H^{*}(BG_{2})\cdot u_{5}$. This completes the proof. ∎
@@ 4.3. The $E_{\infty}$–page and completion of the proof
The restriction to total degree $j\leq 10$ in Theorem 1.1 is motivated by two factors. First, this range is sufficient to capture the first nontrivial cohomological dependence on the bundle class $k$, which manifests via the differential $d_{6}(u_{5})$. Second, extending the analysis beyond degree $10$ requires a substantially more detailed understanding of the fibre cohomology
$M=H^{*}(\Omega^{3}_{0}G_{2};\mathbb{F}_{2})$
as an unstable $\mathcal{A}$–module, which would involve more elaborate Eilenberg–Moore spectral sequence computations beyond the scope of the present paper.
Once $M$ is determined in a larger range, the computation of $H^{*}(B\mathcal{G}_{k};\mathbb{F}_{2})$ should reduce to a finite problem in the “hit” theory over the polynomial algebra $H^{*}(BG_{2})$, and may be amenable to computer implementation. It would also be interesting to compare the present low-degree analysis with Choi’s calculation [5] of the mod $2$ homology of the gauge groups $\mathcal{G}_{k}$ for $G=G_{2}$, via the path-loop fibration of $B\mathcal{G}_{k}$ and the resulting information on $H_{*}(\mathcal{G}_{k};\mathbb{F}_{2})$. We hope to return to these questions in future work.